OpenStudy (anonymous):

evaluate the partial deriv fx(x,y) at the given point f(x,y) =xy ln(y/x)+ln(2x-3y)^2 at p(1,1

OpenStudy (zzr0ck3r):

ok so treat y as a constant and take the derivative with respect to x

OpenStudy (zzr0ck3r):

should that say f(x,y) = ?

OpenStudy (anonymous):

f(x,y)=xyln(y/x)+ln(2x-3y)^2 [p _{o}(1,1)\]

OpenStudy (anonymous):

I'm still a little confused could you do it step by step

OpenStudy (zzr0ck3r):

ok the term on the left xyln(y/x) we treat y as a constant and differentiate with respect to x so use the product rule y*1*ln(y/x) + yx(1/(y/x))*(-y/x^2)

OpenStudy (zzr0ck3r):

this is just the first term, tell me when that makes sense

OpenStudy (zzr0ck3r):

\[y*(x)'*\ln{\frac{y}{x}}+y*x*(\ln{\frac{y}{x}})'=y*1*\ln{\frac{y}{x}}+y*x*\frac{1}{\frac{y}{x}}*(\frac{y}{x})'\]= \[y*1*\ln{\frac{y}{x}}+y*x*\frac{1}{\frac{y}{x}}*\frac{-y}{x^2}\]

OpenStudy (zzr0ck3r):

= \[y*1*\ln{\frac{y}{x}}+y*x*\frac{-1}{x}\]

OpenStudy (zzr0ck3r):

@makylat you left :(

OpenStudy (anonymous):

y∗1∗lnyx+y∗x∗−1x final answer?

OpenStudy (zzr0ck3r):

no, its the first term